OpenStudy (anonymous):

need help asap i have 5 homework to take in 2 hours i dont understand any of it and i need it to pass to next level in school

OpenStudy (anonymous):

. Find the roots of the polynomial equation. x2 – 14x – 32 = 0 (Points : 1) 2 and –16 –6 and 6 –2 and 16 –9 and 8 Question 2. 2. Find the zeros of the polynomial function. f(x) = x2 + 9x + 20 (Points : 1) 2, 10 –4, –5 –2, –10 4, 5 Question 3. 3. Find the zeros of the polynomial function. f(x) = x2 – 4x – 60 (Points : 1) –10, 6 –14, 5 –6, 10 –5, 14 Question 4. 4. Find the roots of the polynomial equation. x2 – 8x – 48 = 0 (Points : 1) –6, 8 –4, 12 –12, 4 –8, 6 Question 5. 5. Find the roots of the polynomial equation. x2 + 24 = –11x (Points : 1) 12, 2 3, 8 –8, –3 –12, –2

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (usukidoll):

wow I think I need to come up with a manual for this as I already explained it in another question... your goal is to look for a specific combination of numbers that can produce the b term and the c term

OpenStudy (usukidoll):

\[x^2 – 14x – 32 = 0\] we need to find what 2 numbers can product 14 and 32. Let's deal with 32.. since there is already multiple choice out there I can rule out all possible combinations and have 16 and 2 Now we just need to find which signs I needs to produce a double negative 16 + 2 = 18.. no... 16-2 = 14 yes! so I have to take these two numbers in reverse 16 as -16 and -2 as 2 in factored form I will have (x-16)(x+2) = 0 solve .. what x's make this equation 0 x=16,-2

OpenStudy (usukidoll):

aww man can product.. can produce x(

OpenStudy (usukidoll):

x^2+2x-16x-32 should give me back x^2-14x-32 so I found my numbers.. can you use this as an example and be able to finish the rest of the problems?

OpenStudy (anonymous):

hate to adment it but im to stupid to solve this by myself. ive been up for two days tryong to finish all my other slasses .-.